OpenStudy (anonymous):

A rectangular box has a square base, and its height is 1 cm longer than the length of each side of its base. If each side of its base has length x cm, show that its total surface area is given by A=6x^2+4x cm^2.

OpenStudy (anonymous):

The two square ends have area x*x=x^2. The other 4 sides have areas x*(x+1)=x^2+x. The total area is therefore 2*x^2+4*(x^2+x)=6x^2+4x

OpenStudy (anonymous):

thank you so much! so to find surface area you have to multiply all the areas of the sides?

OpenStudy (anonymous):

No, you add them - the surface area is the total area of all of the faces. Multiplying them wouldn't give you anything useful.

OpenStudy (anonymous):

And no problem :)